OpenStudy (anonymous):

Prove using coordinate geometry: If a point is on the perpendicular bisector of a segment, then it is equidistant from the endpoints of the segment.

OpenStudy (mertsj):

Could I ask why you reposted this?

OpenStudy (anonymous):

yep i got that what they have given what i have to prove its confusing

OpenStudy (anonymous):

If a point is equidistant from the endpoints of the segment then it is on the perpendicular bisector of the segment. Let [AB] be a given segment. Let M be an arbitrary point in space such that AM = BM. Project M onto (d) and call H the intersection point of (d) and (AB). Consider the triangles AMH and BMH. These two triangles have the angle HBM = the angle HAM (Because the triangle ABM is isosceles in M), and the angle BHM = the angle AHM (= 90 degrees). Therefore, the angle AMH = the angle BMH (two of the angles are the same in both triangles, so the other must be the same) These triangles also have AM = BM, they also have a common side which is [MH]. In summary, we showed that the triangles have two equal sides, and the angles between those sides are also equal. Therefore these triangles are congruent. Therefore AH = BH and thus H is the midpoint of [AB]. Hence, (d) is on the perpendicular bisector of the [AB].

OpenStudy (anonymous):

i gave my teacher this answer and she says its incomplete

OpenStudy (mertsj):

It is not a coordinate geometry proof.

OpenStudy (mertsj):

Coordinate geometry proofs use coordinates to prove them.

OpenStudy (anonymous):

the proof will be completed in two steps. 1) prrove that if P is on the bisector it is equidistant 2) now, proove that if it is equidistant, it HAS to be on the perpendicular bisector!!

OpenStudy (mertsj):

Referring to the diagram I drew previously... the distance from P to A is :

OpenStudy (mertsj):

\[\sqrt{(0--a)^2+(c-0)^2}=\sqrt{a^2+c^2}\]

OpenStudy (mertsj):

Similarly, the distance from P to B is: \[PB=\sqrt{(0-a)^2+(c-0)^2}=\sqrt{a^2+c^2}\]

OpenStudy (anonymous):

brb

OpenStudy (mertsj):

@electrokid Thanks!!

OpenStudy (anonymous):

@Mertsj seems that this question was asked not twice but thrice by the same person on the same day within 20min!!!!

OpenStudy (anonymous):

but the problem is that we have to give reasons regarding each one